OpenStudy (anonymous):

Help!! Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book?

OpenStudy (sweetburger):

Do you know how to start this problem?

OpenStudy (anonymous):

regular price = 18 discount price is 15% of 18.....so for the discount , we turn percent to decimal, then multiply discount = 18 * 0.15 = ??

OpenStudy (sweetburger):

yes correct so then you get the value of 18*.15 then you must take that value and find the amount of tax on the discounted value

OpenStudy (sweetburger):

It would look like this. 18*.15 = x (x = discounted value of the book) so x*.6 = y (y = the sales tax on the book) so your final answer would be x + y

OpenStudy (anonymous):

Its c